### v2

# A Homothetic Gauge Theory

Abstract
While the principle of local gauge invariance forms the bedrock of modern physics, the symmetry of scale invariance has traditionally been treated on a different footing. This paper develops a homothetic gauge theory (HGT) that promotes local scale transformations to a genuine gauge symmetry, placing them within the same geometric framework as conventional interactions. We introduce a novel mathematical structure based on an extended vector bundle that linearizes the action of local scaling, allowing for the definition of a homothetic connection and a corresponding gauge-covariant field strength. A central mathematical innovation of this theory is the development of a homothetic Hodge–de Rham complex and cohomology, wherein the standard differential and codifferential operators are deformed in a scale-covariant manner. When this formalism is applied to classical electromagnetism, as a U(1) gauge theory, the generalized homothetic Maxwell equations are obtained, and a direct connection is established with the spontaneous symmetry breaking. As a significant physical application, this framework provides a compelling resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition, we construct a non-singular electrostatic potential.
Keywords: gauge theory, scale invariance, local symmetries, vector bundles, field theory, Weyl integrable geometry, Hodge–de Rham cohomology, spontaneous symmetry breaking, Maxwell equations, potential theory
classification
1 Introduction
At the heart of modern theoretical physics lies the principle of local gauge invariance, a powerful idea that dictates the fundamental interactions of nature by elevating global symmetries to local, spacetime-dependent ones [8]. One of the most well-established gauge theories in physics is classical electromagnetism, formulated as a $U(1)$ gauge theory. Its mathematical foundation rests on the geometry of principal bundles and connections, offering a prototypical example of how gauge symmetry can encode physical interactions [14]. By extending this geometric framework to incorporate local scale transformations—homotheties—we aim to generalize the notion of gauge invariance itself. Such a generalization offers novel geometric and conceptual reinterpretations of gauge structure, aligning with recent insights into the foundations of gauge symmetries and their reformulations [10].
The paper establishes a new homothetic gauge theory where the gauge freedom corresponds to local changes of scale. The central innovation is the linear structure endowed upon scale transformations, which naturally integrates into the standard fiber bundle formalism. It occurs by embedding the tangent and cotangent bundles in the doubled bundles, anchored by a fixed but arbitrary reference vector field. It transforms the non-linear action of scaling into a linear matrix operation on the fibers of these extended bundles.
This linearization permits the definition of an infinite-dimensional Lie group of fiber-wise homothetic transformations, which serves as the gauge group of the theory. We then proceed to build the essential machinery of a gauge theory upon this foundation: a compatible connection, a covariant derivative that respects the homothetic gauge symmetry, and an associated field-strength (curvature) tensor.
In the paper, the Weyl integrable geometry is obtained as a special case of the general formulation, and it is shown that a homothetic cohomology is definable consistent with this very conformal geometry, resulting in a generalized de Rham complex. When applied to the context of electromagnetism, the theory yields a generalization of Maxwell’s equations. These ”homothetic Maxwell equations” feature new interaction terms coupling the electromagnetic field to the gradient of the scalar gauge field that parameterizes the local scale.
Furthermore, the theory offers a novel and compelling resolution to one of the most persistent problems in classical field theory: the infinite self-energy of a point charge. By formulating the Laplace and Poisson equations within this homothetic framework, we demonstrate that a point source can be modeled as a non-singular object. The divergence is regularized not by altering the fundamental laws of electrodynamics at short distances, but by redefining the nature of a source as a boundary condition on an infinitesimal, penalized interface. This geometric regularization emerges naturally from the theory without ad hoc modifications.
While the use of doubled geometries and extended bundles is not new—notably appearing in string theory (e.g., double field theory) [16] and generalized geometry [13]—our focus on homotheties as a gauge symmetry sets this work apart. Additionally, conformal gauge theories [20] often emphasize angle-preserving transformations, whereas we highlight scale-preserving yet non-isometric maps. The proposed framework thus contributes a unique mathematical structure to the study of spacetime symmetries and potential extensions of general relativity and gauge theory.
The present work was initially began as a penalty method, a mathematical technique aiming to impose the solid boundary on the Navier–Stokes equations via controlling the dynamics of the equations [28, 1]. However, due to its applicability to Maxwell’s equations— as an exact field equation — subsequent geometric interpretations of the method emerged [29]. In fact, penalty methods have been widely used for decades in computational physics as mathematical tools, sometimes accompanied by some physical interpretations [24]. From this perspective, the present theory can be regarded as a geometric interpretation of a particular penalty method for Maxwell’s equations and their subset equations.
A feature of HGT is that it clarifies the usual split between symmetry of the equations and symmetry of their solutions [10, 25]. Instead of purely multiplicative Weyl rescalings, we employ an affine action generated by the dilatation one-form $b=d\lambda$ with offset fields $\phi_{d}$. The field equations remain homothetically covariant, while physically meaningful scales enter through initial/boundary data and gauge choices. In this manner, “spontaneous” breaking of local scale invariance arises from selecting a background/gauge (fixing $\lambda$) and offsets $\phi_{d}$ that act as order parameters, not from a Higgs-type potential. The equations keep homothetic covariance; particular solution classes pick a preferred scale via the imposed data (see $\S$ 7) [33, 25].
The generalized Maxwell sector (Secs. 5.1 &\& 7.1) makes the above issue concrete: the homothetic Lorenz gauge and associated wave/Laplace operators support a penalization that enforces Dirichlet data on a shrinking interface. In the singular limit one recovers the exterior electrovac and avoids the point-charge self-energy divergence; the emergent scale is set by the penalization profile and choice of $\lambda$ (Sec. 8.2), so symmetry is preserved in the equations but broken by solution selection.
An Operational Roadmap of the Paper
We begin by interpreting conformal geometry as a special case of a broader homothetic framework—specifically, as a homothetic geometry with a zero homothety center. This perspective allows us to elevate local scale transformations to a general gauge symmetry. To facilitate this, we introduce a novel mathematical structure based on doubled tangent and cotangent bundles, which linearizes the action of the homothetic group. Upon this foundation, we construct a set of gauge-covariant differential operators, leading to a generalized de Rham complex and an associated Hodge decomposition theorem, which ensures the mathematical consistency of the resulting field theories. As a primary application, we apply this formalism to classical electromagnetism, yielding a set of homothetically covariant Maxwell equations. A distinctive feature of these equations is the emergence of interaction terms that couple the electromagnetic field to the scalar gauge field governing the local scale. These terms persist even under simplifying assumptions and function as geometric penalty terms. This structure allows for the imposition of boundary conditions directly within the field equations, thereby reformulating the theory in a sourceless, singularity-free manner. By imposing a generalized Lorenz gauge condition, we derive homothetic wave and Laplace equations that naturally encode boundary data. This framework is shown to be equivalent to solving classical Dirichlet or Neumann problems via a volumetric penalization method. Finally, we leverage this result to address the classical problem of the infinite self-energy of a point charge. By modeling the source as a penalized Dirichlet boundary condition on a shrinking interface, we construct a non-singular electrostatic potential, resolving the divergence in a manner that emerges naturally from the geometry.
The paper is structured as follows. Section 2 introduces the block–homothetic gauge group and the doubled-bundle construction that renders local scale transformations linear. Section 3 embeds this algebraic setup in a generalized Weyl-integrable geometry, deriving the compatible connection and curvature. In Section 4 we develop the homothetic de Rham complex and prove an extended Hodge decomposition that underpins the subsequent field theory. Section 5 specializes the formalism to electromagnetism, obtaining homothetically covariant Maxwell equations and a modified Lorentz force. Section 6 then derives the associated wave and Laplace equations. Section 7 discusses the relation between HGT and spontaneous symmetry breaking in U(1)-symmetric theories. The physical and mathematical applications of the theory are provided in two related sections. Section 8.1 analyses the Cauchy, Dirichlet, and Neumann boundary-value problems; the discussions that are applied in Section 8.2 to construct a non-singular electrostatic potential for a point charge. Section 9 summarises the main findings. For the sake of continuity, some technical materials are gathered in five appendices at the end of the paper.
2 First setup: Block-homothetic Transformations
From the mathematical viewpoint, we want to have homothetic transformations with linear operations; and it is achievable via embedding the tangent bundles in the doubled spaces. Let $(M,\eta)$ denote $(1+3)$-dimensional Minkowski spacetime, with metric $\eta:=\mathrm{diag}(-,+,+,+)$.
We begin by defining a canonical embedding of tangent and cotangent bundles. For each $x\in M$, consider the tangent space $T_{x}M$. For a fixed, future-directed, timelike vector $X_{d}\in T_{x}M$, we consider a fiber-wise embedding:
$T_{x}M\hookrightarrow T_{x}M\oplus T_{x}M,\qquad X\mapsto\bar{X}=\begin{pmatrix}X\\ X_{d}\end{pmatrix},\qquad\forall X\in T_{x}M.$ (1)
This embedding is well-defined for each $x\in M$ and turns the homothetic transformations to the linear ones in the extended bundle $TM\oplus TM$. Note that since $X_{d}$ is fixed, $\bar{X}$ defines a section in the Whitney-sum bundle $TM\oplus TM$.
Similarly, for the cotangent space, using the musical isomorphism $\flat:TM\to T^{*}M$ induced by $\eta$, given $X^{\flat}_{d}:=\eta(X_{d})$, we define:
$T_{x}^{*}M\hookrightarrow T_{x}^{*}M\oplus T_{x}^{*}M,\qquad X^{\flat}\mapsto\bar{X}^{\flat}:=\begin{pmatrix}X^{\flat}\\ X^{\flat}_{d}\end{pmatrix},\qquad X^{\flat}:=\eta(X)\in T_{x}^{*}M.$ (2)
We now define an infinite-dimensional group of pointwise homothetic transformations acting fiber-wise. Let:
$\mathcal{H}:=\left\{\Lambda[\lambda]\mid\lambda\in C^{\infty}(M,\mathbb{R})\right\},\qquad\Lambda[\lambda](x):=\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\ 0&1\end{pmatrix}$, (3)
where the group operation is given by pointwise matrix multiplication:
$(\Lambda[\lambda_{1}]\Lambda[\lambda_{2}])(x)$ $=\Lambda[\lambda_{1}+\lambda_{2}](x),$
$\Lambda[0](x)$ $=\mathbb{I}_{2}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
As one can see, $\mathcal{H}$ is isomorphic, as a Lie group, to the additive group $C^{\infty}(M,\mathbb{R})$ under pointwise addition.
Remark 1
One can write:
$\Lambda[\lambda(x)]\;=\;e^{-\lambda(x)T}=\;\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\[2.0pt] 0&1\end{pmatrix}$, (4)
where
$T\;=\;\begin{pmatrix}1&-1\\[2.0pt] 0&\,0\end{pmatrix}$. (5)
In the rest of the paper, everywhere that is not confusing, we enjoyed the shorthand writing $\bar{X}\mapsto e^{-\lambda(x)T}\,\bar{X}$, or for the fiberwise operations $X\mapsto e^{-\lambda}\,X$, for $X\in{\mathbb{R}}^{2}$.
The induced fiber-wise homothety on tangent vectors is then:
$\hat{X}:=\begin{pmatrix}\tilde{X}\\ X_{d}\end{pmatrix}:=\Lambda[\lambda]\cdot\bar{X}=\begin{pmatrix}e^{-\lambda(x)}X+\left(1-e^{-\lambda(x)}\right)X_{d}\\ X_{d}\end{pmatrix}$. (6)
This map is well-defined since all components live in the same fiber $T_{x}M$. The first component realizes a homothetic transformation centered at $X_{d}$ with scale $e^{-\lambda(x)}$, while the second component remains fixed; and as one can see, the transformation is linear on the extended space.
Analogously, we define the action of $\mathcal{H}$ on the dual space:
$\hat{X}^{\flat}:=\begin{pmatrix}\tilde{X}^{\flat}\\ X_{d}^{\flat}\end{pmatrix}:=\Lambda[\lambda]\cdot\bar{X}^{\flat}=\begin{pmatrix}e^{-\lambda(x)}X^{\flat}+\left(1-e^{-\lambda(x)}\right)X_{d}^{\flat}\\ X_{d}^{\flat}\end{pmatrix}$. (7)
It should be noted that $\mathcal{H}$ has left actions on both extended fibers $TM\oplus TM$ and $T^{*}M\oplus T^{*}M$.
We then define a diagonal subgroup of $\mathcal{H}\oplus\mathcal{H}$ to capture simultaneous homothetic action on both vectors and covectors:
${\bf H}:=\left\{(\Lambda,\Lambda)\mid\Lambda\in\mathcal{H}\right\}=\Delta(\mathcal{H})\subset\mathcal{H}\oplus\mathcal{H}.$ (8)
Let $\bar{\bar{X}}:=\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}$ and $\hat{\hat{X}}:=\begin{pmatrix}\hat{X}\\ \hat{X}^{\flat}\end{pmatrix}$. Then:
$\hat{\hat{X}}=\begin{pmatrix}\Lambda&0\\ 0&\Lambda\end{pmatrix}\begin{pmatrix}\bar{X}\\ \bar{X}^{\flat}\end{pmatrix}={\bf H}\cdot\bar{\bar{X}}.$ (9)
We now equip the extended bundle $TM\oplus TM$ with the metric $\bar{\eta}:=\eta\oplus\eta$. This allows us to extend the musical isomorphism fiber-wise:
$\bar{\flat}_{x}:T_{x}M\oplus T_{x}M\to T_{x}^{*}M\oplus T_{x}^{*}M,\qquad\bar{X}\mapsto\bar{X}^{\bar{\flat}}:=(\eta(X),\eta(X_{d}))=(X^{\flat},X_{d}^{\flat}).$ (10)
This will play a central role in the homothetic Hodge decomposition developed in section 4.
[…]
Figure 1: Fiberwise embedding of the tangent space $T_{x}M$ into the extended space $T_{x}M\oplus T_{x}M$ using a fixed vector $X_{d}$. This facilitates linear realization of homothetic transformations.
Finally, for general differential forms $\phi,\phi_{d}\in\Omega^{k}(M)$ with $0\leq k\leq 4$, where $\phi_{d}$ is fixed, we define the embedding:
$\Omega^{k}(M)\hookrightarrow\Omega^{k}(M)\oplus\Omega^{k}(M),\qquad\phi\mapsto\bar{\phi}=\begin{pmatrix}\phi\\ \phi_{d}\end{pmatrix},\qquad\forall\phi\in\Omega(M).$ (11)
And $\mathcal{H}$ has a well-defined pointwise action on $\bar{\phi}$:
$\hat{\phi}:=\Lambda\cdot\bar{\phi}=\begin{pmatrix}\tilde{\phi}\\ \phi_{d}\end{pmatrix}=\begin{pmatrix}e^{-\lambda(x)}\phi+\left(1-e^{-\lambda(x)}\right)\phi_{d}\\ \phi_{d}\end{pmatrix}$. (12)
This construction extends naturally to bundles of differential forms, allowing us to analyze gauge-like homotheties and their induced effects on geometric structures such as the Hodge decomposition.
The above construction, while self-contained, is not arbitrary. As we demonstrate in Appendix A, this entire framework arises naturally as a specific gauge fixing within the standard formulation of conformal Cartan geometry, providing a deeper geometric justification for our approach. Moreover, in appendix C the connection and curvature of $\mathcal{H}$ bundle, as a principal bundle, are obtained.
3 A Generalized Weyl Integrable Geometry
On the extended space $\bar{\bar{X}}$, a symmetry manifests itself: the transformation $\bar{\bar{X}}\to\hat{\hat{X}}={\bf H}\bar{\bar{X}}$ preserves the musical isomorphism $\bar{\flat}$. In particular, it is due to the commutativity of $\bar{\flat}$ and $\Lambda$:
$\Lambda\bar{X}^{\bar{\flat}}=\Lambda\bar{\eta}\bar{X}=\bar{\eta}(\Lambda\bar{X})=(\Lambda\bar{X})^{\bar{\flat}}.$ (13)
However, this does not imply that the lengths and angles of vectors are preserved under the transformation $\bar{\bar{X}}\mapsto\hat{\hat{X}}$—only that they remain unchanged within a local Lorentz frame. To make this distinction precise, we define an affine symmetric 2-form $\tilde{\eta}:TM\times TM\to{\mathbb{R}}$ via:
[…]
where $\tilde{X}$ and $\tilde{Y}$ are given by the homothetic action in Eq. (6). A direct computation yields:
$\tilde{\eta}(X,Y)=e^{-2\lambda}\eta(X,Y)+(1-e^{-\lambda})^{2}\eta(X_{d},X_{d})+e^{-\lambda}(1-e^{-\lambda})[\eta(X,X_{d})+\eta(Y,X_{d})].$ (15)
As one can see, this affine metric kernel reflects explicitly the existence of the offset vector $X_{d}$ and the scaling $\lambda(x)$. It is worth mentioning that for $\lambda=0$, it reduces to $\tilde{\eta}=\eta$, and for $\lambda\to\infty$, it degenerates to:
$\tilde{\eta}(X,Y)\to\eta(X_{d},X_{d}),$ (16)
independently of $X$ and $Y$. The structure encoded in Eq. (15) thus reflects the symmetry (13) that naturally combines a local scaling $\lambda(x)$ with a fiberwise translation defined by $X_{d}(x)$, consistent with the action of the group $\mathcal{H}$.
While this offset structure is geometrically natural and consistent with our next cohomological framework, however, the general 2-form (15), with $X_{d}\neq 0$, is hard to justify as a physical spacetime metric. It arises from the linearization of the homothetic map, and it is the most direct consequence of our initial setup. However, it is not a metric in the usual sense. In particular, the presence of $X_{d}$ in $\tilde{\eta}$ means that even the zero vector $X=0$ can acquire a nontrivial norm— violates the usual requirement that $\|0\|=0$. Therefore, to recover a physical spacetime geometry that supports unambiguous geodesic separation, we impose a physicality condition by restricting ourselves to the subspace where the offset (homothety center) vector vanishes:
$X_{d}=0,$ (17)
which leads to the conventional conformal geometry:
$\boxed{\tilde{\eta}(X,Y)=e^{-2\lambda(x)}\eta(X,Y).}$ (18)
That is, to this point, the conformal geometry is interpreted as a homothetic geometry with zero homothety center.
It should be stressed that although we finally set $X_{d}=0$, this is not just a complicated way to achieve the standard conformal geometry. In fact, we will always consider this geometry as a homothetic geometry with the zero homothety center. In particular, it manifests itself in the differential forms we will associate with it in the next section— see Eq. (20) and the discussions after that.
On the other hand, nothing in our formulation prevents us from admitting the conventional Weyl structure. Recall that [30] a Weyl integrable geometry is defined by the triple $(M,[\eta],\phi)$, where $\eta]$ is the conformal class of Lorentzian metrics and $\phi:=d\lambda$ is a globally defined, closed 1-form satisfying
$\nabla_{\rho}\eta_{\mu\nu}=-2\phi_{\rho}\eta_{\mu\nu}.$ (19)
This non-metricity condition defines a Weyl-compatible connection that preserves null directions but allows local rescaling governed by $\phi$. However, in our formulation, both the conformal metric and the 1-form $\phi:=d\lambda$ arise naturally from the homothetic structure. In particular, the non-metricity 1-form $\phi=d\lambda$ appears automatically from the homothetic scaling parameter.
4 Homothetic Hodge Decomposition
In this section, the Hodge-de Rham theory is established for the homothetically transformed differential forms, defined on the above conformally transformed spacetime.
4.1 The Homothetically Transformed Differential Forms
Given the differential forms $\phi,\phi_{d}\in\Omega^{\bullet}(M)$, where $\phi_{d}$ is fixed, we define the homothetically transformed differential form $\tilde{\alpha}$ as:
$\tilde{\alpha}:=e^{-\lambda(x)}\alpha+\left(1-e^{-\lambda(x)}\right)\alpha_{d},$ (20)
obtainable from the first row of $\hat{\alpha}=\Lambda\bar{\alpha}$. Associating such a transformation to the conformally transformed spacetime is mathematically legitimated (see appendix A and the discussion below). Moreover, the following special cases are noticeable:
  • For $\alpha_{d}=0$, it reduces to the familiar Weyl-type scaling $\tilde{\alpha}=e^{-W\lambda(x)}\alpha$, with the Weyl weight $W=1$. In contrast to $X_{d}$, here we will keep $\alpha_{d}\neq 0$ in general.
  • A homothetically invariant form (equivalent to $W=0$ in the classical Weyl geometry) is obtainable by setting:
    $\alpha=\alpha_{d},$ (21)
    that is, $\lambda(x)$ does not affect $\alpha$ because:
    $\tilde{\alpha}=e^{-\lambda(x)}(\alpha-\alpha_{d})+\alpha_{d}=e^{-\lambda(x)}(\alpha_{d}-\alpha_{d})+\alpha_{d}=\alpha_{d}.$ (22)
    As will be seen, even for the homothetically invariant forms (such as electromagnetic gauge potentials, as discussed in $\S$ 5), the general homothetic formulation has still new and interesting things.
The physical motivation and mathematical justification of the homothetic transformation (20) are in order.
Physical motivation
Consider the limiting case $\lambda(x)\to\infty$, where the conformal factor approaches zero, $e^{-2\lambda(x)}\to 0$, and the associated volume form vanishes, $(d\text{vol}_{\tilde{\eta}})_{x}\to 0$. A theory that remains valid across all scales requires that the fields themselves converge to a well-defined state in this limit. We postulate that for any initial form $\alpha(x)$, the transformed field $\tilde{\alpha}(x)$ must approach a universal asymptotic form $\alpha_{d}(x)$:
$\lim_{\lambda\to\infty}\tilde{\alpha}(x)=\alpha_{d}(x).$
This principle endows the offset field $\alpha_{d}$ with a clear physical interpretation: it is the ground state to which all other forms are driven as the local spacetime metric degenerates. The existence of such a non-trivial asymptotic field is a profound consequence of the homothetic framework.
Mathematical formalism
This physical intuition can be framed rigorously using the language of contraction mappings on the cotangent bundle. At each point $x\in M$, the homothetic transformation (20) defines a fiberwise map:
$\mathbf{T}_{\lambda(x)}:T_{x}^{*}M\to T_{x}^{*}M,\qquad\alpha\mapsto\mathbf{T}_{\lambda(x)}(\alpha):=e^{-\lambda(x)}\alpha+\left(1-e^{-\lambda(x)}\right)\alpha_{d}.$ (23)
We endow each fiber $T_{x}^{*}M$ with the norm $\|\cdot\|_{x}$ induced by the background Minkowski metric $\eta$. For any two forms $\alpha_{1},\alpha_{2}\in T_{x}^{*}M$, the distance between their images is:
$\|\mathbf{T}_{\lambda(x)}(\alpha_{1})-\mathbf{T}_{\lambda(x)}(\alpha_{2})\|_{x}=\|e^{-\lambda(x)}(\alpha_{1}-\alpha_{2})\|_{x}=e^{-\lambda(x)}\|\alpha_{1}-\alpha_{2}\|_{x}.$
For any $\lambda(x)>0$, the Lipschitz constant $e^{-\lambda(x)}$ is strictly less than 1, meaning $\mathbf{T}_{\lambda(x)}$ is a strict contraction on the complete metric space $(T_{x}^{*}M,\|\cdot\|_{x})$. By the Banach Fixed-Point Theorem, this map has a unique fixed point. Solving for the fixed point $\alpha_{fp}$ gives:
$\alpha_{fp}=e^{-\lambda(x)}\alpha_{fp}+(1-e^{-\lambda(x)})\alpha_{d}\quad\implies\quad\alpha_{fp}=\alpha_{d}.$
Thus, for any positive scaling $\lambda(x)$, the offset form $\alpha_{d}$ is the unique, stable equilibrium point of the transformation in each fiber.
This construction defines a smooth family of fiberwise contractions across the manifold $M$, forming a pointwise contraction semigroup. As the scaling becomes infinite, the map converges to a projection that sends every form to this unique fixed point:
$\lim_{\lambda(x)\to\infty}\mathbf{T}_{\lambda(x)}(\alpha)=\alpha_{d},\quad\forall\alpha\in T_{x}^{*}M.$
This provides a rigorous mathematical justification for the physical principle that the theory must remain well-defined in the limit of a vanishing local metric. The implications of this contraction become central to our generalized Hodge theory.
How these transformations propagate through the Hodge-de Rham complex is the question that will be addressed in the next subsection, after developing some operators we need for their analysis.
4.2 Generalized de Rham complex and Hodge decomposition
The key step in dealing with the homothetically transformed differential forms is the appropriate construction of the differential operators on this geometry. In this regard, we first define:
$\bar{d}:=\begin{pmatrix}d&0\\ 0&d\end{pmatrix},\quad\bar{\delta}:=\begin{pmatrix}\delta&0\\ 0&\delta\end{pmatrix},\quad\bar{\Delta}:=\begin{pmatrix}\Delta&0\\ 0&\Delta\end{pmatrix},\qquad\ \bar{\star}:=\begin{pmatrix}\star&0\\ 0&\star\end{pmatrix}$$\quad$ (24)
where $d$ is the exterior derivative, $\delta$ is the codifferential, $\Delta=d\delta+\delta d$ is the Hodge–de Rham Laplacian, and $\star$ is the Hodge star operator. Then, we define these operators on the homothetically transformed space of differential forms as:
Generalized exterior derivative:
$\hat{d}:=\begin{pmatrix}d+(d\lambda)\wedge&-(d\lambda)\wedge\\ 0&d\end{pmatrix}$. (25)
Generalized codifferential:
$\hat{\delta}:=\begin{pmatrix}\delta-*(d\lambda)\wedge*&*(d\lambda)\wedge*\\ 0&\delta\end{pmatrix}$. (26)
Generalized Laplacian:
$\hat{\Delta}:=\hat{\delta}\hat{d}+\hat{d}\hat{\delta}$. (27)
Block Hodge star operator:
$\hat{\star}$$\bar{\star}$. (28)
The key features of these operators are that (it can be verified directly):
$\boxed{\hat{d}\Lambda=\Lambda\bar{d},\qquad\hat{\delta}\Lambda=\Lambda\bar{\delta},\qquad\hat{\Delta}\Lambda=\Lambda\bar{\Delta},\qquad\bar{\star}\Lambda=\Lambda\bar{\star},}$ (29)
and as the direct consequences of these features, one can verify that irrespective of the rank:
  1. 1.
    Both $\hat{d}$ and $\hat{\delta}$ are nilpotent, that is, $\hat{d}^{2}=0$ and $\hat{\delta}^{2}=0$.
  2. 2.
    With the conformally scaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$, $\hat{\delta}$ is the adjoint of $\hat{d}$  [^2] It should be noted that without the conformal scaling $\eta\to\tilde{\eta}$, $\hat{\delta}$ is not the formal adjoint of $\hat{d}$. It may be seen as an evidence of internal consistency of the theory., that is:
    $\langle\hat{d}\hat{\alpha},\hat{\beta}\rangle_{\tilde{\eta}}=\langle\hat{\alpha},\hat{\delta}\hat{\beta}\rangle_{\tilde{\eta}}$ (30)
  3. 3.
    As a consequence of adjointness of $\hat{d}$ and $\hat{\delta}$, one can say that $\hat{\Delta}$ is elliptic.
  4. 4.
    The block Hodge star operator $\bar{\star}$ and $\Lambda$ commute without any further modification.
Geometric and Physical Interpretations
The generalized operators $\hat{d},\hat{\delta},\hat{\Delta}$, $\hat{\star}$ are not arbitrary: they arise naturally from the demand that the differential structure be compatible with the homothetic transformations introduced earlier. Specifically, the exterior derivative and codifferential must transform covariantly under the homothetic action $\Lambda$, preserving the structure of the deformed forms $\hat{\phi}=\Lambda\bar{\phi}$. This compatibility is encoded in the identities (29) which imply that homothetic transformations commute with the differential operators, effectively rendering the homothetic structure gauge covariant.
Physically, this suggests that the transformation group $\mathcal{H}$ plays a role analogous to a gauge group, albeit non-compact and infinite-dimensional. The field $\lambda(x)$ behaves like a scalar gauge parameter—more precisely, a logarithmic scale factor—whose differential $d\lambda$ enters the transformation of both $d$ and $\delta$ in a manner reminiscent of a connection 1-form in Weyl geometry. The antisymmetric contributions involving $d\lambda\wedge$ and $d\lambda\wedge*$ indicate that this is a kind of abelian connection with scale-coupled torsion.
Moreover, the ellipticity of $\hat{\Delta}$ ensures that the generalized Laplacian still governs the dynamics of field propagation and harmonic analysis on the manifold. From a gauge-theoretic point of view, this confirms that no essential degrees of freedom are lost through the transformation: instead, they are reorganized in a manner compatible with the homothetic group action.
[…]
Figure 2: Diagrammatic representation of the generalized Hodge-de Rham complex with homothetic differential forms $\hat{\Omega}^{k}$, showing both the generalized exterior derivative $\hat{d}$ and codifferential $\hat{\delta}$. At each degree $k$, the kernel of the generalized Laplacian $\hat{\Delta}$ identifies the harmonic representatives $\mathcal{H}^{k}$ in the generalized Hodge decomposition $\mathcal{H}^{k}:=\ker\hat{\Delta}^{k}$ (${\mathcal{H}}^{k}$ not to be confused with the homothety group $\mathcal{H}$).
Cohomological Consequences
Importantly, the nilpotency of $\hat{d}$ allows us to define a generalized de Rham cohomology:
$\hat{H}^{k}(M):=\ker(\hat{d}^{k})/\text{Im}(\hat{d}^{k-1}),$
which encodes the topological information of the manifold as filtered through the homothetic structure. While the underlying topology of $M$ remains unchanged, the representatives of each cohomology class are now deformed by the scalar field $\lambda(x)$ and the offset $\alpha_{d}$. When $\lambda=0$, all modifications vanish and we recover standard de Rham theory. However, for $\lambda\neq 0$, the cohomology classes are parametrized by homothetically equivalent forms.
This leads us to the generalized Hodge decomposition:
$\boxed{\hat{\Omega}^{k}(M)=\text{Im}~\hat{d}^{k-1}~\oplus~\text{ker}~\hat{\Delta}^{k}~\oplus~\text{Im}~\hat{\delta}^{k+1},}$ (31)
in which the space of homothetically transformed $k$-forms splits orthogonally into exact, coexact, and harmonic components, all defined relative to the Weyl-rescaled metric $\tilde{\eta}=e^{-2\lambda(x)}\eta$. This decomposition underlies the forthcoming formulation of field equations—most notably, generalized Maxwell-type theories—in which the harmonic representatives encode gauge-invariant field configurations within the homothetic background. A functional analytic proof of decomposition (31) is provided in appendix D.
5 Homothetic Electromagnetism
The above homothetic Hodge-de Rham complex can be constructed for the electromagnetism, as a $U(1)$ gauge theory defined on the Minkowski spacetime manifold. It means embedding any $k$-form $\alpha^{k}$, where $0\leq k\leq 4$ in $\Omega^{k}(M)\oplus\Omega^{k}(M)$, where $\mathcal{H}$ can act on, making $\hat{\Omega}^{k}$. However, in order for the homothetic transformations to preserve the symmetries of the electromagnetism, they must act directly on the connections of the principal bundle, that is, the gauge potential fields $A=A_{\mu}dx^{\mu}$ (as the $\mathfrak{u}(1)$-valued 1-forms on $M$). Such a setup results in Fig. 3, in which we defined:
$\bar{A}:=\begin{pmatrix}A\\ A_{d}\end{pmatrix},\qquad\hat{A}:=\begin{pmatrix}\tilde{A}\\ A_{d}\end{pmatrix}=\Lambda\bar{A},$ (32)
and the other $k\neq 1$ forms are transformed with the offset forms $\delta A_{d}$, $dA_{d}$, $d^{2}A_{d}=0$, etc., inducing by this genuine transformation. For example, by definition of:
$\bar{F}:=\begin{pmatrix}F\\ F_{d}\end{pmatrix},\qquad\hat{F}:=\begin{pmatrix}\tilde{F}\\ F_{d}\end{pmatrix}=\Lambda\bar{F},$ (33)
the offset is $F_{d}:=dA_{d}$, where $A_{d}$ has already defined.
5.1 Sourceless Maxwell Equations
On the 4-dimensional Minkowski spacetime $(M,\eta)$, classical (source-free) electromagnetism is expressed as:
$dF=0,$ (34)
$\delta F=0,$ (35)
with $F=dA$ for some 1-form $A$ defined on $M$. We now consider an offset potential $A_{d}\in\Omega^{1}(M)$ with $F_{d}=dA_{d}$, and define the block forms $\hat{A}$ and $\hat{F}$ as above, and the Maxwell equations in the homothetically transformed geometry obtains as:
$\hat{d}\hat{F}=0,$ (36)
$\hat{\delta}\hat{F}=0.$ (37)
Then, $\hat{F}=\hat{d}\hat{A}$ is invariant under the homothetic gauge transformation:
$\boxed{\hat{A}\mapsto\hat{A}+\hat{d}\hat{\phi}}\qquad\text{where}\qquad\hat{\phi}=\Lambda\bar{\phi}=\Lambda\begin{pmatrix}\phi\\ \phi_{d}\end{pmatrix}$, (38)
and $\phi,\phi_{d}$ are arbitrary smooth functions on $M$. This is the homothetic equivalent of the classical gauge invariance of the Maxwell equations.
The first row of the homothetic Maxwell equations (36)-(37) yields:
$d\tilde{F}+(d\lambda)\wedge(\tilde{F}-F_{d})=0,$ (39)
$\delta\tilde{F}-*(d\lambda)\wedge*(\tilde{F}-F_{d})=0.$ (40)
As one can see, some additional terms have appeared which represent the interaction of the electromagnetic fields with the offset fields. Apparently, for $d\lambda\equiv 0$, these additional terms disappear, and the classical Maxwell equations retrieve.
[…]
Figure 3: Homothetic Hodge-de Rham complex for electromagnetism (the lower part), and its relation with the classical complex (the upper part). The homothetic transformation genuinely occurs on the connection 1-forms, and it induces other transformations on the other $k\neq 1$ differential forms.
By translation of the above equations on the tangent bundle, the generalized Maxwell equations are obtained in the more conventional three-vector formulation:
$\nabla\cdot\tilde{E}+\nabla\lambda\cdot(\tilde{E}-E_{d})$ $=0,\qquad\text{generalized Gauss Law};$ (41)
$\partial_{t}\tilde{E}-\nabla\times\tilde{B}+\partial_{t}\lambda(\tilde{E}-E_{d})-\nabla\lambda\times(\tilde{B}-B_{d})$ $=0,\qquad\text{generalized Ampère's Law};$ (42)
$\partial_{t}\tilde{B}+\nabla\times\tilde{E}+\partial_{t}\lambda(\tilde{B}-B_{d})+\nabla\lambda\times(\tilde{E}-E_{d})$ $=0,\qquad\text{generalized Faraday Law};$ (43)
$\nabla\cdot\tilde{B}=0,\quad\text{for }\nabla\cdot B_{d}$ $=0,\qquad\text{Gauss's Law for Magnetism}.$ (44)
This formulation offers a homothetically covariant generalization of electromagnetism. A distinguishing feature of this system is that it is sourceless. In fact, the emergence of penalty terms, with geometric origins, enable us to derive the dynamics of electromagnetic fields without the need for source terms. This issue is crucial in construction of a non-singular Model for a Point Electric Charge in $\S$ 8.2. However, for the sake of completeness, a non-homogeneous formulation is also provided in $\S$ 7.
6 The Homothetic Lorenz gauge, Wave, and Laplace Equations
We now turn to the derivation of the dynamical equations that arise from homothetically transformed electromagnetic potentials and scalar fields. These include a wave equation and a Laplace equation defined on the homothetically transformed geometry.
6.1 Homothetic Harmonicity of the Gauge Potential
Starting from the homothetic Maxwell equations (36)-(37), and using $\hat{F}=\hat{d}\hat{A}$, we choose the generalized Lorenz gauge:
$\boxed{\hat{\delta}\hat{A}=0.}$ (45)
This leads to the homothetic harmonic equation:
$\hat{\Delta}\hat{A}=0.$
Hence, the transformed potential $\hat{A}$ satisfies a generalized wave equation governed by $\hat{\Delta}$.
Now, let the block scalar potential be:
$\hat{\phi}:=\begin{pmatrix}\tilde{\phi}\\ \phi_{d}\end{pmatrix}$,
with $A=d\phi$, and $A_{d}=d\phi_{d}$ by similar reasoning as for $\hat{A}$, one obtains:
$\hat{\Delta}\hat{\phi}=0.$
Then, from the generalized Lorenz gauge (45), with:
$\hat{A}=\begin{pmatrix}\tilde{A}\\ A_{d}\end{pmatrix}$,
and the form of $\hat{\delta}$, one obtains:
$\delta\tilde{A}-*(d\lambda)\wedge*(\tilde{A}-A_{d})=0.$
This yields a generalized conservation-like equation:
$\boxed{\nabla\cdot\tilde{A}-\nabla\lambda\cdot(\tilde{A}-A_{d})=0.}$ (46)
In spacetime index notation:
$\partial_{\mu}\tilde{A}^{\mu}-\partial_{\mu}\lambda(\tilde{A}^{\mu}-A_{d}^{\mu})=0,$
which modifies the classical condition $\partial_{\mu}A^{\mu}=0$ by a geometric source term tied to $\lambda$.
6.2 Wave and Laplace Equations
We now derive the homothetic wave and Laplace equations for $\tilde{\phi}$, the first component of $\hat{\phi}$. Recall that $\hat{\Delta}\hat{\phi}=0$, and write the d’Alembertian operator as:
$\square:=\partial_{\mu}\partial^{\mu}=-\partial_{t}^{2}+\nabla^{2},\quad|\partial\lambda|^{2}:=\partial_{\mu}\lambda\,\partial^{\mu}\lambda,\quad(\partial\lambda\cdot\partial f):=\partial_{\mu}\lambda\,\partial^{\mu}f.$
General Case
$\boxed{\square\tilde{\phi}+2\,\partial_{\mu}\lambda(\partial^{\mu}\tilde{\phi}-\partial^{\mu}\phi_{d})+(\square\lambda+|\partial\lambda|^{2})(\tilde{\phi}-\phi_{d})=0}$
Stationary Fields
Assuming $\lambda$ and the potentials are time-independent:
$\boxed{\nabla^{2}\tilde{\phi}+2\,\nabla\lambda\cdot(\nabla\tilde{\phi}-\nabla\phi_{d})+(\nabla^{2}\lambda+|\nabla\lambda|^{2})(\tilde{\phi}-\phi_{d})=0}$ (47)
As one can see, these equations exhibit new coupling between the scalar potential components $\tilde{\phi}$ and $\phi_{d}$, mediated by $\lambda$.
7 Mathematics or Physics— HGT and the Spontaneous Symmetry Breaking
The principle of local scale invariance is a long-standing concept in theoretical physics. In Weyl’s original theory, fields transform multiplicatively. HGT proposes a different, affine-style transformation: $\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d})$. This introduces a fixed-point offset field, $\phi_{d}$, fundamentally altering the symmetry’s structure. This section argues that demanding invariance under this homothetic action is a form of spontaneous symmetry breaking (SSB). Meanwhile, the dynamics of $\lambda$ and the offset fields $\phi_{d}$ remain as the open questions.
First, it is important to recall two logically distinct types of scale transformations. In the classical Weyl approach [25], one considers local rescalings of the metric and fields as:
$g_{\mu\nu}\mapsto e^{2\sigma(x)}g_{\mu\nu},\qquad\phi\mapsto e^{W\sigma(x)}\phi,$ (48)
where $W$ is the Weyl weight of the field $\phi$. A non-Weyl-invariant model can be made invariant by introducing a compensating scalar field $\tau$ that transforms as $\tau\mapsto\tau-\sigma$. By dressing other fields with powers of $e^{\tau}$, one can achieve invariance under Weyl transformations. This is commonly interpreted as a spurionic or Stückelberg restoration of symmetry. Fixing a gauge (or assigning a vacuum expectation value to $\tau$) then breaks the symmetry spontaneously and realizes it nonlinearly. [^3] We ignore quantum anomaly issues here; the discussion is entirely classical.
In HGT, an analogous procedure applies, but the transformations act differently:
$g_{\mu\nu}\mapsto e^{-2\lambda(x)}g_{\mu\nu},\qquad\phi\mapsto\phi_{d}+e^{-\lambda(x)}(\phi-\phi_{d}),$ (49)
where $\lambda$ is the logarithm of the local scale factor, and $\phi_{d}$ is an offset field obeying the same equations as $\phi$ and transforming in the same representation. By defining $\delta\phi:=\phi-\phi_{d}$, the transformation simplifies to $\delta\phi\mapsto e^{-\lambda}\delta\phi$, meaning that $\delta\phi$ behaves like a conventional weight 1-1 field under local scalings. This encapsulates the essence of HGT transformations and allows the action to retain form-invariance under local homotheties.
Spontaneous symmetry breaking [11, 33] occurs when the underlying equations or action of a theory respect a symmetry, but the chosen vacuum state does not. In gauge theories, this typically manifests through the introduction of a scalar field acquiring a nonzero vacuum expectation value (VEV), thus selecting a specific vacuum and reducing the full symmetry to a residual subgroup. Importantly, SSB does not eliminate the symmetry but conceals it: physical configurations transform nontrivially under the broken generators. In classical field theory, this principle underlies both the Higgs mechanism and geometric symmetry reductions. Within the context of HGT, SSB becomes the bridge between mathematical symmetry and emergent physical structure, as is exemplified by the generalized Maxwell equations in the sequel.
In $\S$ 5.1, we derived the source-free generalized Maxwell equations directly from the homothetic Hodge–de Rham theory. This geometry-based approach establishes the mathematical foundation of the equations. Regardless of their physical interpretation, the homothetic gauge freedom and the derived equations are of interest in their own right, with potential applications to the theory of partial differential equations and to computational physics. Here we shall derive the same equations, in the presence of the current density $J\neq 0$, from a process that closely reflects spontaneous symmetry breaking.
The starting point is an action that is not manifestly scale-invariant due to explicit dependence on a metric and field configurations. By introducing the homothetic variables $(\phi,\phi_{d})$ and employing the transformation rule in Eq. (49), the action is modified in such a way that symmetry is effectively restored at the level of its form. This mirrors the logic of Stückelberg restoration: one modifies the theory so that it is invariant under a larger symmetry group, and the breaking is realized through a fixed gauge or background choice.
This viewpoint implies that HGT treats the effect of local scale transformation as a spontaneous symmetry breaking mechanism. However, unlike traditional gauge theories with Higgs-like scalars, the dynamical origins and physical interpretations of the fields $\lambda$ and $\phi_{d}$ remain unresolved. The field $\lambda(x)$ governs local rescalings but is not yet derived from a fundamental action. Likewise, the offset field $\phi_{d}$ is introduced to enable covariant transformations, but its dynamics and role in observable quantities have not been clarified. These gaps must be addressed if HGT is to evolve into a physical rather than purely formal framework.
7.1 Example: The Generalized Maxwell Equations and Scale-symmetry Restoration
For our derivations, we recall the following relations (from $\S$ 4.2):
  1. 1.
    The block differential on $k$-forms,
    $\hat{d}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}\ :=\ \begin{pmatrix}d\alpha+b\wedge\alpha-b\wedge\beta\\[2.0pt] d\beta\end{pmatrix},\qquad b:=d\lambda,$ (50)
    and the adjoint $\hat{\delta}:=-\hat{*}\hat{d}\hat{*}$ with respect to $\tilde{\eta}$, where $\hat{*}$ is the Hodge operator on the block bundle built from $_{\tilde{\eta}}$ on each component.
  2. 2.
    The block field-strength $\hat{F}:=\hat{d}\,\hat{A}$ for a one-form doublet $\hat{A}:=\big{(}\tilde{A},A_{d}\big{)}^{\mathsf{T}}$, with upper component
    $\tilde{F}:=d\tilde{A}+b\wedge(\tilde{A}-A_{d}),\qquad F_{d}:=dA_{d}.$ (51)
Start from the standard Maxwell system on $(M,\eta)$ coupled to a conserved current $J$,
$S_{\rm Max}[A;J]\ =\ -\frac{1}{2}\int F\wedge*_{\eta}F\ +\ \int A\wedge*_{\eta}J,\qquad F:=dA,\quad d*_{\eta}J=0,$ (52)
which in $d=4$ is conformally invariant in the sourceless sector but, in the presence of sources and a fixed choice of scale, need not be Weyl-invariant in the sense of (48). To restore local scale symmetry homothetically, introduce the compensator $\lambda$, the offset potential $A_{d}$, and the doublet $\hat{A}=(\tilde{A},A_{d})^{\mathsf{T}}$ with
$\tilde{A}\ :=\ A_{d}+e^{-\lambda}(A-A_{d}),\qquad\text{so that}\qquad\tilde{A}-A_{d}=e^{-\lambda}(A-A_{d}).$ (53)
We then define the HGT Maxwell action by
$S_{\rm HGT}[\,\hat{A};\lambda;\hat{J}\,]\ :=\ -\frac{1}{2}\int\big{\langle}\hat{F},\hat{F}\big{\rangle}_{\tilde{\eta}}\ +\ \int\big{\langle}\hat{A},\hat{J}\big{\rangle}_{\tilde{\eta}},\qquad\hat{F}:=\hat{d}\,\hat{A}$, (54)
where $\hat{J}:=(\tilde{J},J_{d})^{\mathsf{T}}$ is a source doublet (the top slot couples to the physical sector; the bottom slot is available for regularizations/defects and can be set to zero in the simplest applications). By construction, $S_{\rm HGT}$ is exactly invariant under local homotheties (49) with $b=d\lambda$, since all derivatives are the Weyl-covariant ones and the pairing uses $\tilde{\eta}$.
Varying $\hat{A}$ yields the generalized Maxwell equations
$\boxed{\quad\hat{d}\,\hat{F}\ =\ 0,\qquad\hat{\delta}\,\hat{F}\ =\ \hat{J},\quad}$ (55)
i.e. the Weyl-covariant Bianchi identity and source equation on the block bundle. In a Weyl-covariant Lorenz gauge,
$\hat{\delta}\,\hat{A}=0\quad\Longleftrightarrow\quad\tilde{\nabla}_{\mu}\tilde{A}^{\mu}\;-\;b_{\mu}\big{(}\tilde{A}^{\mu}-A_{d}^{\mu}\big{)}\ =\ 0,$ (56)
the equations reduce to a block wave equation
$\hat{\Delta}\,\hat{A}\ =\ \hat{J},\qquad\hat{\Delta}:=\hat{d}\,\hat{\delta}+\hat{\delta}\,\hat{d}$, (57)
with $\hat{\Delta}$ the HGT Laplace–de Rham operator built from $\tilde{\eta}$ and $b=d\lambda$. Writing out only the physical (upper) slot, one obtains the compact form
$\delta_{\tilde{\eta}}\!\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ \tilde{J},\qquad d\left[d\tilde{A}+b\wedge(\tilde{A}-A_{d})\right]\ =\ 0,$ (58)
where $\delta_{\tilde{\eta}}$ is the codifferential with respect to $\tilde{\eta}=e^{-2\lambda}\eta$. When $A_{d}$ is chosen as a pure gauge or fixed reference (e.g. $A_{d}\equiv 0$) these reduce to
$\delta_{\tilde{\eta}}\!\left(d\tilde{A}+b\wedge\tilde{A}\right)=\tilde{J},\qquad d\left(d\tilde{A}+b\wedge\tilde{A}\right)=0,$ (59)
which are the Maxwell equations in an integrable Weyl geometry with scale connection $b=d\lambda$. Equations (56)–(59) make manifest that the only effect of the restoration is to (i) replace ordinary by Weyl-covariant differentials and (ii) introduce a mixing with the offset field $A_{d}$ through $\tilde{A}-A_{d}$, i.e. through the homogeneously transforming difference.
8 Applications
From here on, some applications of the theory will be discussed. Two related mathematically and physically important problems will be tackled in the framework of the homothetic gauge theory, and some new results will be provided and discussed.
8.1 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
According to the left part of Fig. 4, let $S\subset\mathbb{R}^{d}$ be a closed, smooth, orientable, embedded (d1)(d-1)-dimensional submanifold, where $d\in\{1,2,3\}$. By the Jordan–Brouwer separation theorem, $\mathbb{R}^{d}\setminus S$ has exactly two connected components: one bounded (denote it by $\Omega_{i}$) and one unbounded (denote it by $\Omega_{o}$). The closures $\overline{\Omega_{i}}$ and $\overline{\Omega_{o}}$ are smooth $d$-dimensional manifolds with boundary, sharing the boundary $\partial\overline{\Omega_{i}}=\partial\overline{\Omega_{o}}=S$. Then, we consider $S_{i}$ and $S_{o}$ as the interior and exterior sides of $S$, respectively.
Note 1
From here on, we provide a general formulation, which is valid equally for both the interior manifold $\overline{\Omega_{i}}$ and the exterior manifold $\overline{\Omega_{o}}$. Therefore, we avoided repeatedly using of indices ”ii” and ”oo”.
We denote by $\delta_{n}$  [^4] Not to be confused with the codifferential in the previous sections a sequence of regularizations of the Dirac delta function $\delta_{S}$ (it can be $S_{i}$, or $S_{o}$), such that $\lim_{n\to\infty}\delta_{n}=\delta_{S}$. Then, in the generalized Laplace equation (47), we choose (see Appendix B for justification):
$\lambda_{n}=\ln(1+\delta_{n}),$ (60)
resulting in:
$\nabla^{2}\phi_{n}+2\,\frac{\nabla\delta_{n}}{1+\delta_{n}}\cdot(\nabla\phi_{n}-\nabla\phi_{d})+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0.$ (61)
We observe that, in the singular limit $n\to\infty$, the regularization $\delta_{n}$ approaches the Dirac distribution $\delta_{S}$, supported on the neighborhood of the hypersurface $S$. In this limit, the terms involving derivatives of $\delta_{n}$ become highly localized near $S$, and act to impose both Dirichlet and Neumann conditions simultaneously on the surface.
As a result, equation (61) defines two decoupled Cauchy problems in $\mathbb{R}^{d}$: one on $\overline{\Omega_{i}}$ and one on $\overline{\Omega_{o}}$. Unless one solution is the harmonic continuation of the other, a discontinuity (in the solution or its gradient) appears at $S$. If only a Dirichlet condition $\phi_{S}=\phi_{d}$ is imposed, a discontinuity in the normal derivative $\nabla\phi_{S}\cdot\mathbf{n}$ may appear across $S$. Conversely, if only the Neumann condition $\nabla\phi_{S}\cdot\mathbf{n}=\nabla\phi_{d}\cdot\mathbf{n}$ is enforced, a jump in the solution may appear at $S$. These phenomena are closely related to classical potential theory, particularly the behavior of single-layer and double-layer potentials [21, 22].
Specifically, a single-layer potential yields a continuous potential across the interface but typically introduces a discontinuity in its normal derivative. In contrast, a double-layer potential maintains continuity of the normal derivative but allows for a jump in the potential itself [5]. The behavior of equation (61) in the limit $n\to\infty$ is thus analogous to a combined layer potential, where both Dirichlet and Neumann data are enforced.
${\mathbb{R}}^{d}$$\mathbf{\hat{n}}_{o}$$\mathbf{\hat{n}}_{i}$$\Omega_{i}$$\Omega_{o}$$S_{o}$$S_{i}$
[…]
Figure 4: Left: The closures $\overline{\Omega_{o}}$ and $\overline{\Omega_{o}}$ are considered as manifolds with boundary, sharing the boundary $S$. Right: A manifold with boundary is locally isomorphic (the isomorphism is denoted by $\psi$) with a half space ${\mathbb{H}}^{d}$.
Interestingly, equation (61) also provides a systematic framework for solving classical boundary value problems. Suppose one intends to solve a pure Dirichlet problem—either on the interior or exterior of $S$. In this case, one may first omit the gradient penalizing terms and consider the equation:
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0.$ (62)
This corresponds to a standard Laplace equation perturbed by a sharply localized penalty term, which enforces $\phi_{n}\approx\phi_{d}$ on $S$ in the limit $n\to\infty$. As a result, a Dirichlet boundary condition is imposed through a volumetric penalization, a method reminiscent of embedded boundary techniques and diffuse interface methods [26, 24].
Once $\phi_{n}$ is obtained from equation (62), the normal derivative $\nabla\phi_{n}\cdot\mathbf{n}$ on $S$ is obtained as well. In the context of elliptic boundary value problems, this is equivalent to applying the Dirichlet-to-Neumann (DtN) map $\Lambda$ (not to be confused with the homothetic transformation matrix in the previous sections), often denoted by
$\Lambda:\phi_{n}|_{S}\mapsto\nabla\phi_{n}\cdot\mathbf{n}|_{S}.$ (63)
Now, by assuming $\nabla\phi_{d}:=\nabla\phi_{n}$, the term $(\nabla\phi_{n}-\nabla\phi_{d})$ in the original equation (61) vanishes identically. Thus, one has in effect solved the complete equation (61) under a self-consistent assumption that the constructed $\phi_{d}$ is aligned with the solution gradient on $S$.
A similar argument holds for solving Neumann problems. In this case, one may omit the Dirichlet penalty term and retain only the gradient penalty term:
$\nabla^{2}\phi_{n}+2\,\frac{\nabla\delta_{n}}{1+\delta_{n}}\cdot(\nabla\phi_{n}-\nabla\phi_{d})=0.$ (64)
This imposes a Neumann boundary condition by penalizing the difference between the gradient and the prescribed one $\nabla\phi_{d}$. Upon obtaining a solution $\phi_{n}$, one can define $\phi_{d}:=\phi_{n}$ on $S$, and again observe that the residual term $(\phi_{n}-\phi_{d})$ in equation (61) vanishes. Hence, the full equation is satisfied again under the assumption of gradient-data consistency.
This selective process regularizes the non-regular Cauchy problem and simplifies numerical treatment. It also clarifies the role of the full equation as a unified interface formulation, capable of encoding a range of classical boundary conditions. Moreover, the correspondence with the DtN operator can also be interpreted as a feature of the interface as a generalized boundary operator in the sense of nonlocal elliptic theory [6, 27].
8.2 A Non-Singular Model for a Point Electric Charge
The classical model of electrostatics, governed by the Poisson equation with a delta function source, has long faced a well-known difficulty: the infinite energy of the vacuum associated with point charges [17, 19]. Specifically, when modeling a point charge located at the origin, the potential in $\mathbb{R}^{3}$ takes the form
$\phi(r)=\frac{C}{r},$
where $C$ is a constant and $r=\|\mathbf{x}\|$. While this solution satisfies the Laplace equation away from the origin, it becomes singular at $r=0$, and the corresponding electrostatic energy,
$E=\frac{1}{2}\int_{\mathbb{R}^{3}}|\nabla\phi|^{2}\,dV,$
diverges due to the integrable singularity at the origin. This divergence has been a foundational issue in classical field theory, prompting numerous attempts at regularization, ranging from distributional reformulations [12, 4] to the introduction of extended charge models [7, 2]. In the framework developed in the previous section, a new perspective emerges.
We consider a spherical interface $\partial B(R)\subset\mathbb{R}^{3}$, and solve the generalized Laplace equation with Dirichlet-type penalization:
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0,$ (62 revisited)
where $\delta_{n}$ is a regularization of the Dirac delta function supported on the surface of a ball, $\delta_{\partial B(R)}$, and $\phi_{d}$ is the prescribed Dirichlet potential
$\phi_{d}=\frac{C}{R},$
enforcing a Coulomb-like condition at $r=R$ .
Remark 2
It should be emphasized that the delta function is defined on the surface of the ball $\partial B(R)$, not at the origin. It makes an essentially different situation in comparison with a point delta function. A hollow sphere $S^{2}$ and a point are homotopically different objects irrespective of $R$, and they have different capacitance (from the potential theory viewpoint) irrespective of $R$.
Due to the highly localized nature of the penalty term, the solution $\phi_{n}$ adjusts sharply near the interface to match the boundary value, and decays as $\phi_{n}(r)\sim\frac{C}{r}$ for large $r$, consistent with the behavior of the classical potential in the far field.
[…]
Figure 5: A non-singular point charge model with a spherical interface at $r=R$. The delta function is defined at the surface of the sphere. The interior potential is decoupled from the exterior, where the solution decays as $\phi_{n}\sim\frac{C}{r}$.
The structure of the equation ensures a natural decoupling of the interior and exterior domains. In the limit $n\to\infty$, the penalization enforces the boundary condition only on $\partial B(R)$, and the interior and exterior solutions become independent. Thus, the behavior of the solution in $\mathbb{R}^{3}\setminus B(R)$ is determined entirely by the enforced Dirichlet data and is unaffected by the field inside the ball.
Remark 3
It should be emphasized that we are not aware of the solution inside the ball $B(R)$. Inside the ball is another manifold in which the solution is determined from the boundary condition on its boundary.
We can now safely approach $R\to 0$, by which the ball $B(R)$ gets smaller, and the prescribed boundary potential $\phi_{d}=C/R$ grows, although neither $\partial B(R)$ changes to a point nor $\phi_{d}$ gets infinity. Moreover, since our domain is the exterior domain, and the potential inside the ball plays no role in the exterior solution, the energy,
$E=\frac{1}{2}\int_{\mathbb{R}^{3}\setminus B(R)}|\nabla\phi_{n}|^{2}\,dV,$ (65)
remains finite in the limit $R\to 0$, as the core is effectively excluded from the domain of integration.
Therefore, we have a physically faithful, Coulombian potential associated with a point-like source, but without the divergence of vacuum energy. The source is interpreted not as a singular point but as the limit of a penalized interface of vanishing size, carrying an enforced boundary potential. The infinite energy problem is thereby resolved through a combination of geometrically localized penalization and the decoupling of interior and exterior fields.
This construction should be contrasted with other historical approaches to resolving the divergence of self-energy. Many attempts, such as non-linear electrodynamics theories like the Born-Infeld model [3, 9], modify the fundamental dynamics of the electromagnetic field everywhere in spacetime to regularize the field at the origin. Our approach is fundamentally different. It is a form of geometric regularization that does not alter the linear Maxwell equations in the vacuum (where $\lambda=0$).
9 Conclusions
In this work, we have developed a Homothetic Gauge Theory (HGT) that systematically promotes local scale invariance to the status of a fundamental gauge principle, placing it on the same geometric footing as the interactions of the Standard Model. The central challenge of incorporating the non-linear operation of scaling into the linear framework of gauge theory was overcome by introducing a novel fiber-wise embedding into a doubled vector bundle. This construction linearizes the homothetic transformations, allowing them to be represented by a simple matrix group acting on the extended fibers.
Upon this foundation, we constructed the full machinery of a consistent gauge theory. A generalized Weyl-integrable geometry emerged, in which the spacetime metric is conformally rescaled by a dynamical scalar gauge field, $\lambda(x)$. The mathematical integrity of the theory was secured by developing a homothetic Hodge–de Rham complex, complete with generalized differential operators and a corresponding Hodge decomposition theorem. This framework ensures that the resulting field equations are well-posed and their solutions are unique under appropriate boundary conditions.
The primary physical application of HGT was demonstrated in the context of classical electromagnetism. The theory yields a set of homothetically covariant Maxwell equations that are inherently sourceless. Traditional source terms are replaced by geometric penalty terms that couple the electromagnetic field to the gradient of the scale field, $\nabla\lambda$. This unique feature allows boundary conditions to be encoded directly into the dynamics, reformulating boundary-value problems as sourceless field equations over the entire domain.
It has shown that local scale covariance is preserved by the HGT field equations while preferred scales arise from the selection of gauge/background data. In particular, the dilatation one–form $b=d\lambda$ and the offset fields $\phi_{d}$ act as order parameters: fixing $\lambda$ and $\phi_{d}$ selects solution sectors that break scale symmetry without introducing a Higgs-type potential. Thus, HGT realizes spontaneous symmetry breaking (SSB) as a property of solutions (and their data), not of the equations themselves.
Within homothetic electromagnetism (Secs. 56), the SSB viewpoint explains how physical scales enter the generalized Maxwell theory in a controlled way. The penalization framework and homothetic Lorenz gauge make the mechanism explicit: the penalization profile and the chosen $\lambda$ determine the emergent length scale, while the equations remain homothetically covariant. This solution-level SSB suggests several directions: coupling the mechanism to gravity (Cartan/Weyl-integrable interpretations), analyzing anomalies and quantization of the homothetic sector, studying dynamical selection of $\lambda$ and $\phi_{d}$ in matter-coupled models, and exploring additional interface limits beyond electrostatics. We expect these lines to sharpen the relation between mathematical scale covariance and the emergence of physical scales across HGT applications.
From the computational physicist’s viewpoint, perhaps the most significant physical consequence of this framework is a natural resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source via a volume penalization method that enforces a Dirichlet condition in the limit, we constructed a non-singular electrostatic potential that is finite everywhere, including at the origin. The divergence is eliminated not by an ad-hoc modification of Maxwell’s laws, but as a direct consequence of the underlying geometry. This result suggests that what we perceive as sources may be geometric interfaces where the scale of spacetime undergoes rapid change.
Finally, by showing that the entire formalism can be understood as a specific gauge fixing within the broader context of conformal Cartan geometry, we have grounded HGT in a well-established mathematical tradition. The theory presented here is not merely a mathematical reformulation but a predictive framework with the potential for broad application. Future work could involve extending this formalism to non-Abelian gauge groups or exploring its implications for gravity, where the homothetic gauge field might offer new insights into dark energy or the cosmological constant problem.
Appendix A Cartan Geometric Interpretation of the Gauge Structure
In this appendix we reformulate the block–homothetic gauge group of Section 2 within the standard framework of conformal Cartan geometry. Besides making contact with familiar conventions in the literature, this embedding singles out the fixed vector $X_{d}$ used in the main text and shows that every field that appears there is the section of an associated bundle.
A.1 Conformal Cartan Geometry
Let $G=\mathrm{SO}(4,2)$ be the four-dimensional conformal group and $P\subset G$ its parabolic subgroup that stabilises a null ray. A conformal Cartan geometry on a smooth spacetime $M$ is a principal $P$-bundle $\pi:\mathcal{P}\to M$ endowed with a $\mathfrak{g}$-valued 1-form
$A\;\in\;\Omega^{1}\!\bigl{(}\mathcal{P};\,\mathfrak{g}\bigr{)}$
that is (i) an isomorphism $T_{p}\mathcal{P}\xrightarrow{\;\simeq\;}\mathfrak{g}$ at every $p\in\mathcal{P}$, (ii) $P$-equivariant $R_{a}^{*}A=\operatorname{Ad}(a^{-1})A$, and (iii) reproduces the generators of the right action: $A(\zeta^{*})=\zeta$ for all $\zeta\in\mathfrak{p}$.
The $|1|$-grading $\mathfrak{g}=\mathfrak{p}_{-1}\oplus\mathfrak{p}_{0}\oplus\mathfrak{p}_{1}$ decomposes the Cartan connection into
$A=e^{\mu}P_{\mu}+\omega^{\mu\nu}M_{\mu\nu}+b\,D+f^{\mu}K_{\mu},$
where $e^{\mu}$ is the solder form, $\omega^{\mu\nu}$ the spin connection, $b$ the Weyl (dilational) gauge field, and $f^{\mu}$ the special-conformal gauge field.
Gauge fixing to the block–homothetic subgroup.
From now on we freeze a future-directed timelike (equivalently : a null) ray in $\mathbb{R}^{4,2}$. On the base $M$ that choice projects to the fixed vector $X_{d}\!\in\!TM$ introduced in Section 2. The remaining structure group is the semidirect product $H=\mathrm{SO}(3,1)\ltimes\mathbb{R}$; its $\mathbb{R}$ factor acts by the block-upper-triangular matrices
$\Lambda[\lambda](x)=\begin{pmatrix}e^{-\lambda(x)}&1-e^{-\lambda(x)}\\[2.0pt] 0&1\end{pmatrix},\qquad\lambda\in C^{\infty}(M).$ (66)
On a weight-zero tractor $V=\bigl{(}V~~V_{d}\bigr{)}^{\!\top}$ this reads
$\Lambda[\lambda](x)\cdot V=\begin{pmatrix}e^{-\lambda(x)}\,V+\bigl{(}1-e^{-\lambda(x)}\bigr{)}V_{d}\\[2.0pt] V_{d}\end{pmatrix}$, (67)
which reproduces Eq. (6) when we put $V\mapsto X$ and $V_{d}\mapsto X_{d}$.
Associated bundles and differential forms.
Because every representation of $P$ (and hence of $H\subset P$) yields an associated bundle, all fields used in the main text acquire a natural Cartan meaning:
  • vector fields of homothetic weight $w$ are sections of $\mathcal{P}\times_{P}\mathbb{R}^{1,3}[w]$;
  • $p$-forms of weight $w$ are sections of $\mathcal{P}\times_{P}\!\bigl{(}\Lambda^{p}\mathfrak{p}_{-1}^{*}\bigr{)}[w]$;
  • the doubled bundle $TM\oplus TM$ of Section 2 is the direct sum of two tractor bundles carrying weights 0 and 1-1, respectively.
Hence all operators in $\S$ 2 are functorial in Cartan geometry.
A.2 Spontaneous Symmetry Breaking and Scale Selection
Cartan geometry also explains how the conformal symmetry is broken in stages:
  1. 1.
    The full conformal group $G=\mathrm{SO}(4,2)$ breaks to the Poincaré subgroup $H=\mathrm{ISO}(3,1)$ once the solder form $e^{\mu}$ and spin connection $\omega^{\mu\nu}$ acquire non-degenerate vacuum values.
  2. 2.
    A further breaking $H\to\mathrm{SO}(3,1)$ occurs after a local scale is chosen. In our formulation that choice is equivalent to fixing the vector $X_{d}$ or, dually, to giving $b$ an expectation value.
[…]
Figure 6: Symmetry-breaking chain induced by the solder form and the scale selector $X_{d}$ (or, equivalently, the Weyl field $b$).
A.3 Interpretation of the Fixed Vector $X_{d}$
The canonical embedding $X\mapsto\bar{X}=\bigl{(}X~~X_{d}\bigr{)}^{\!\top}$ used in the main text corresponds, in the Cartan picture, to the choice of a weight 1-1 compensator tractor $X_{d}$. Its presence linearizes local scale transformations, exactly as a compensator field does in Weyl gravity.
In summary, the block–homothetic group of Section 2 is nothing but the residual $H$-structure that survives the simple gauge fixing (66). All fields and operations of the main text extend functorially to tractors, so the entire theory fits consistently inside conformal Cartan geometry.
Appendix B Finding $\lambda$ for imposition of boundary conditions
Here, in the absence of a dynamical theory for $\lambda(x)$, we fix it in a way that allows us to apply boundary conditions. The result has the potential for applications in computational physics. The method presented here is not a complete solution to the mathematical problem of finding $\lambda$, but the $\lambda$ found is acceptable for sufficiently regular boundaries.
We employ the well-established theory of regular singular points in the theory of second-order ordinary differential equations (ODEs). Consider the one-dimensional generalized Laplace equation (47):
$\phi^{\prime\prime}+2\lambda^{\prime}(\phi^{\prime}-\phi_{d}^{\prime})+(\lambda^{\prime\prime}+(\lambda^{\prime})^{2})(\phi-\phi_{d})=0,$ (68)
where $(\cdot^{\prime})$ means derivative with respect to $x\in{\mathbb{R}}$, $\phi_{d}:{\mathbb{R}}\to{\mathbb{R}}$ is a given smooth function, and we want to solve the above equation to find $\phi(x)$.
Now, we want the solution to be localized in the neighborhood of the boundary (here the origin $x=0$). Therefore, the coefficients $\lambda^{\prime}$ and $\lambda^{\prime\prime}+(\lambda^{\prime})^{2}$ must go to infinity in the neighborhood of the boundary $x=0$; and it makes $x=0$ a singular point of the equation. Then, to be $x=0$ a regular singular point, we must have:
$\lambda^{\prime}\sim\frac{1}{x}\qquad\lambda^{\prime\prime}+(\lambda^{\prime})^{2}\sim\frac{1}{x^{2}}.$ (69)
Both these conditions satisfy by assuming:
$\boxed{\lambda=\ln(1+\delta_{n}),}$ (70)
where $\delta_{n}$ is a regularization of the Dirac delta function, with the regularization parameter $n$.
By substitution of this $\lambda$ in the equation, one obtains:
$\phi^{\prime\prime}+2\frac{\delta_{n}^{\prime}}{1+\delta_{n}}(\phi^{\prime}-\phi_{d}^{\prime})+\frac{\delta_{n}^{\prime\prime}}{1+\delta_{n}}(\phi-\phi_{d})=0.$ (71)
And since $\delta_{n}\sim\frac{1}{x}$ in the vicinity of $x=0$, one can verify that:
$\frac{\delta_{n}^{\prime}}{1+\delta_{n}}\sim\frac{1}{x},\qquad\frac{\delta_{n}^{\prime\prime}}{1+\delta_{n}}\sim\frac{1}{x^{2}},\qquad\text{in the neighborhood of}~x=0.$ (72)
For the multi-dimensional problems one can use the multi-dimensional Dirac delta function and the appropriate distance functions.
Physical vs. coordinate singularities for $\lambda=\ln\!\bigl{(}1+\delta_{n}\bigr{)}$
In a Weyl integrable spacetime $(M,g,\lambda)$ any two frames $(g,\lambda)$ and $(\bar{g}=e^{2\sigma}g,\;\bar{\lambda}=\lambda-\sigma)$ are physically equivalent. Hence a vanishing conformal factor $\Omega:=e^{-\lambda}$ per se does not signal a spacetime singularity; one must inspect gauge–invariant curvature scalars [^5] Typical choices are $e^{\lambda}R$, $e^{2\lambda}R_{\mu\nu}R^{\mu\nu}$, $e^{2\lambda}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, all invariant under Weyl rescalings. and geodesic completeness.
Now, based on above rationales, let
$\lambda(r)=\ln\!\bigl{[}1+\delta_{n}(r)\bigr{]},\qquad\delta_{n}\in\mathcal{C}^{\infty}_{0}(\mathbb{R}^{3}),$
where $\delta_{n}$ is a regularized Dirac delta, that is, a smooth, compactly–supported mollifier for the Dirac delta. Because $\delta_{n}$ and all of its derivatives are bounded, so are $\lambda$, $\nabla\lambda$ and $\Box\lambda$. For any background metric $g_{\mu\nu}$ that is regular at $x_{0}$ we have (e.g. in four dimensions)
$e^{\lambda}R\;=\;6\!\left[(\nabla\lambda)^{2}-\Box\lambda\right],\qquad e^{2\lambda}R_{\mu\nu}R^{\mu\nu}=\tfrac{3}{2}\!\bigl{[}(\nabla\lambda)^{2}-\Box\lambda\bigr{]}^{2},$
and analogous expressions for higher invariants. All of them remain finite because the square brackets are bounded. Furthermore, the affine–parameter integral along any causal geodesic, $\int\Omega^{2}\,ds,$ diverges only if $|\lambda|\to\infty$ faster than $\ln s$, which never happens for the present $\lambda$.
Conclusion.
The would–be “zero” of $\Omega=e^{-\lambda}$ at any $x_{0}$ inside the support of $\delta_{n}$ is therefore a coordinate artefact: all curvature scalars are finite and causal geodesics can be extended through $x_{0}$. The conformal factor regularises the monopole–like $1/r$ behavior without introducing a physical singularity.
Appendix C Principal $\mathcal{H}$-bundle formulation and curvature
Let $\mathcal{H}\cong({\mathbb{R}}_{>0},\cdot)$ be the homothetic group acting on each doubled fibre by block-diagonal multiplication. [^6] Locally—on a chart (U,x)(U,x)—a homothety is specified by a smooth function $\lambda\in C^{\infty}(U)$, the action being $v\mapsto e^{\lambda(x)}v$. Pointwise multiplication turns $\mathcal{H}$ into a $1$-dimensional abelian Lie group with Lie algebra $\mathfrak{h}\cong{\mathbb{R}}$ under addition. Choose a good open cover $\{U_{\alpha}\}_{\alpha\in A}$ of $M$ and define transition functions
$g_{\alpha\beta}\colon U_{\alpha}\cap U_{\beta}\;\longrightarrow\;\mathcal{H},\qquad g_{\alpha\beta}(x)\;=\;e^{\lambda_{\alpha\beta}(x)},\quad\lambda_{\alpha\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta}),$
satisfying the usual cocycle conditions $g_{\alpha\alpha}=1,\;g_{\alpha\beta}g_{\beta\gamma}=g_{\alpha\gamma}$. They define a principal $\mathcal{H}$-bundle  $\pi\colon P\to M$ with local trivializations $\psi_{\alpha}:\pi^{-1}(U_{\alpha})\!\to\!U_{\alpha}\times\mathcal{H}$ and right action $(p,h)\!\mapsto\!ph$. Throughout we fix a principal connection
$\omega\;\in\;\Omega^{1}(P,\mathfrak{h}),\qquad R\,=\,d\omega\;\in\;\Omega^{2}(P,\mathfrak{h}),$
characterised by:
$\text{\emph{(i)}}\;\;\;R_{X^{\#}}\omega=1\quad\text{and}\quad\text{\emph{(ii)}}\;\;\;R_{h}^{*}\omega=Ad_{h^{-1}}\!\omega\;(=\omega\text{ since }\mathcal{H}\text{ is abelian}),$
with $X^{\#}$ the fundamental field generated by $X\in\mathfrak{h}$.
Local expression.
Choosing local sections $\sigma_{\alpha}\colon U_{\alpha}\to P$ we obtain the connection $1$-forms
$\Gamma_{\alpha}\;:=\;\sigma_{\alpha}^{\!*}\omega\;\in\;\Omega^{1}(U_{\alpha},\mathfrak{h})\quad\text{and}\quad R_{\alpha}\;:=\;\sigma_{\alpha}^{\!*}R\;=\;d\Gamma_{\alpha}.$
Under a change of trivialization one has
$\Gamma_{\beta}\;=\;g_{\alpha\beta}^{-1}\Gamma_{\alpha}\,g_{\alpha\beta}+g_{\alpha\beta}^{-1}\,dg_{\alpha\beta}\;=\;\Gamma_{\alpha}+d\lambda_{\alpha\beta},$ (73)
because $Ad_{g_{\alpha\beta}^{-1}}$ is trivial. Curvature transforms accordingly as $R_{\beta}=R_{\alpha}$, so $R$ is gauge-invariant. If we write $\Gamma_{\alpha}=d\lambda_{\alpha}$ for a smooth scalar $\lambda_{\alpha}\in C^{\infty}(U_{\alpha})$, then (73) is consistent with $\lambda_{\beta}=\lambda_{\alpha}+\lambda_{\alpha\beta}$. Hence $R=d^{2}\lambda_{\alpha}=0$ locally, and globally
$R\;=\;d\Gamma+\Gamma\wedge\Gamma\quad(\text{Maurer–Cartan form}),$
establishing Eq. (3.6) in a manifestly principal-bundle language. The global existence of $\Gamma$ is equivalent to the exactness of the Čech $1$-cocycle $\{\lambda_{\alpha\beta}\}$ in $H^{1}(M,R)$.
Appendix D A functional-analytic proof for the homothetic Hodge decomposition (31)
Throughout this appendix $(M,g)$ denotes a compact oriented $n$-manifold without boundary; the case with boundary is analogous once standard Dirichlet or Neumann data are imposed. Fix a smooth homothetic gauge potential $\lambda\in C^{\infty}(M)$ and abbreviate
$d_{\lambda}\;:=\;e^{-\lambda}d\,e^{\lambda},\qquad\delta_{\lambda}\;:=\;e^{\lambda}\delta\,e^{-\lambda},\qquad\Delta_{\lambda}\;:=\;\delta_{\lambda}d_{\lambda}+d_{\lambda}\delta_{\lambda}.$
We adopt Sobolev norms $\|\cdot\|_{H^{s}}$ based on the Levi-Civita connection; all statements are independent of that choice.
Lemma 1 (Ellipticity)
The operator $\Delta_{\lambda}:\Omega^{k}(M)\to\Omega^{k}(M)$ is a second-order strongly elliptic differential operator with the same principal symbol as the ordinary Laplace–de Rham operator $\Delta_{g}$.
Proof 1
Write $d_{\lambda}=d+(\,d\lambda\wedge\cdot\,)$; the added zeroth-order term does not affect the principal symbol, hence $\sigma_{2}(\Delta_{\lambda})=\sigma_{2}(\Delta_{g})$. Strong ellipticity follows because $\Delta_{g}$ is strongly elliptic (positive-definite symbol) and lower-order terms preserve that property [32].
Lemma 2 (Gårding–type estimate)
For every $s\in{\mathbb{N}}$ there exists $C_{s}>0$ such that
$\|\xi\|_{H^{s+2}}\;\leq\;C_{s}\bigl{(}\,\|\Delta_{\lambda}\xi\|_{H^{s}}+\|\xi\|_{H^{s}}\bigr{)},\quad\forall\,\xi\in C^{\infty}\Omega^{\bullet}(M).$
Proof 2
Combine ellipticity (Lemma 1) with an Agmon–Douglis–Nirenberg parametrix construction [15]. The lower-order terms introduced by $\lambda$ change only the constant in the estimate.
Corollary 1 (Fredholmness)
$\Delta_{\lambda}:H^{s+2}\!\to\!H^{s}$ is Fredholm with index 0 for every $s\in{\mathbb{N}}$.
Proof 3
Lemma 2 implies compact embedding of the unit ball of $H^{s+2}$ in $H^{s+1}$, whence Rellich compactness yields a parametrix $P:H^{s}\!\to\!H^{s+2}$ modulo compact terms. Fredholm index is stable under compact perturbations and $\ker\Delta_{\lambda}$ is finite-dimensional.
Proof 4 (Homothetic Hodge Decomposition (31))
Equip [^7] Because $e^{\pm\lambda}$ are bounded on compact $M$, $\langle\!\cdot,\cdot\!\rangle_{\lambda}$ is equivalent to the usual $L^{2}$ inner product. $\Omega^{\bullet}(M)$ with the inner product $\langle\!\alpha,\beta\!\rangle_{\lambda}=\!\int_{M}\alpha\wedge*e^{-2\lambda}\!\beta$ and observe $\delta_{\lambda}$ is the formal adjoint of $d_{\lambda}$. By Cor. 1, $\Delta_{\lambda}$ extends to a self-adjoint unbounded operator in $L^{2}$ with compact resolvent; hence its spectrum is discrete and every $\xi\in L^{2}\Omega^{\bullet}$ decomposes orthogonally as
$\xi\;=\;d_{\lambda}\alpha\;+\;\delta_{\lambda}\beta\;+\;\gamma,\quad\gamma\in\ker\Delta_{\lambda},$
where $\alpha,\beta$ are unique up to the kernel of $\delta_{\lambda}$ and $d_{\lambda}$, respectively. Regularity theory implies $\alpha,\beta,\gamma$ are smooth whenever $\xi$ is, completing the Hodge decomposition. Harmonic forms $\gamma$ constitute a finite-dimensional vector space isomorphic to the de Rham cohomology $H^{\bullet}(M,R)$ because $d_{\lambda}$ is cohomologous to $d$. $\square$
Remark 4
If $\partial M\neq\varnothing$, impose absolute ($\imath_{\nu}\xi=0,\;\imath_{\nu}d_{\lambda}\xi=0$) or relative ($\xi|_{\partial M}=0,\;d_{\lambda}\xi|_{\partial M}=0$) boundary conditions; Lemma 2 adapts following [31] and the decomposition remains valid.